### v1

# Emergent Matryoshka doll-like point gap in a non-Hermitian quasiperiodic lattice

Abstract
We propose a geometric series modulated non-Hermitian quasiperiodic lattice model, and explore its localization and topological properties. The results show that with the ever-increasing summation terms of the geometric series, multiple mobility edges and non-Hermitian point gaps with high winding number can be induced in the system. The point gap spectrum of the system has a Matryoshka doll-like structure in the complex plane, resulting in a high winding number. In addition, we analyze the limit case of summation of infinite terms. The results show that the mobility edges merge together as only one mobility edge when summation terms are pushed to the limit. Meanwhile, the corresponding point gaps are merged into a ring with winding number equal to one. Through Avila’s global theory, we give an analytical expression for mobility edges in the limit of infinite summation, reconfirming that mobility edges and point gaps do merge and will result in a winding number that is indeed equal to one.
preprint: APS/123-QEDpreprint: This line only printed with preprint option
I Introduction
Anderson localization reveals the phenomenon of how electron Bloch wave diffusion is suppressed by random disorder PWAnderson1958 . To this day, the discussion of localization problems remains a dynamic field in condensed matter physics. The scaling theory established in the middle of last century fully reveals that Anderson localization caused by disorder is dependent on system dimension EAbrahams1979 ; PALee1985 ; FEvers2008 ; BHetenyi2021 . Specifically, in low dimensions ($D<3$), arbitrarily small disorder suffices to localize the system EAbrahams1979 . For the case of $D=3$, extended and localized states can coexist, separated by the so called mobility edge (ME) FEvers2008 ; ALagendijk2009 ; NFMott1967 ; JKoo1975 . This implies that, the localization phase transition depends on the disorder strength and energy in 3D system.
Different from random disorder, quasiperiodic systems, which are intermediate between disorder and order system, are widely used to study localization phase transitions and MEs in low-dimensional systems. A widely known $1D$ quasiperiodic system is called the Aubry-André (AA) model, whose transition point from extended to localized phases can be exactly determined by self-duality conditions SAubry1980 ; SYJitomirskaya1999 . The self-duality condition also predicts that the extended and localized states cannot coexist and that there are no MEs. Consequently, generalized versions of the AA model have been proposed to study MEs, including modulated lattice structures, modified quasiperiodic potentials APadhan2022 ; SDasSarma1988 ; SDasSarma1990 ; TLiu2022 ; SGaneshan2015 ; XLi2017 ; HYao2019 ; XLi2020 ; YCZhang2022 ; XCZhou2023 ; YWang2020 ; XPLi2016 ; ZLu2022 ; SRoy2021 , and the introduction of long-range hopping JBiddle2011 ; XXia2022 ; XDeng2019 ; MGoncalves2023 ; JBiddle2010 . Beyond these, the AA model has found widespread applications in exploring topological Anderson insulators JLi2009 ; GQZhang2021 ; DWZhang2020b ; DWZhang2020c , transport properties YEKraus2012 ; JCCharlier2007 ; SHuang2024 , entanglement transitions LZhou2024 ; LZhou2024a ; LZhou2023 ; SZLi2024 , and multifractal phases XBeng2019 ; MSarkar2022 ; SZLi2023 ; XLin2023 ; HYao2019 ; MGoncalves2023 ; FLiu2015 ; XCZhou2023 . These remarkable properties can be experimentally realized using cold atomic systems FAAn2018 ; DTanese2014 ; GModugno2010 ; MLohse2016 ; SNakajima2016 ; DWZhang2018 , optical lattices FAAn2018 ; HPLuschen2018 ; SAGredeskul1989 ; DNChristodoulides2003 ; YEKraus2012 ; TPertsch2004 ; HPL2018 ; MSchreiber2015 ; MLohse2016 ; SNakajima2016 , photonic crystals YLahini2009 ; MVerbin2013 ; MVerbi2015 ; PRoushan2017 ; PRoushan2017 , and superconducting qubits HLi2023 .
Recently, there has been a growing interest in localization and topological phase transitions in non-Hermitian quasiperiodic systems PHe2022 ; SZLi2024b ; SZLi2024a ; APadhan2024 ; DWZhang2020b ; HJiang2019 ; SLonghi2019 ; YLiu2020 ; QBZeng2020a ; TLiu2020 ; XCai2021 ; YLiu2021a ; JClaes2021 ; LZTang2021 ; SLonghi2021a ; SLonghi2021b ; LZhou2022 ; XCai2022 ; WHan2022 ; QLin2022 ; YLiu2021 ; LJZhai2020 ; LZTang2022 ; TQian2024 ; YLiu2021c ; QBZeng2020b , where non-Hermiticity can stem from non-reciprocal hopping or complex on-site potential SLonghi2019 ; HJiang2019 ; QBZeng2020b . Essentially, the two types of non-Hermitian modulations are related. One can uncover this relation through the dual transformation, that complex on-site potential is equivalent to non-reciprocal hopping in the dual space, vice versa. Furthermore, when there are MEs in the system, the above two non-Hermitian modulation will exhibit the dual relationship. Specifically, under the condition of complex on-site potential (non-reciprocal hopping) modulation, the extended states (localized states) possess the real spectra, while the localized states (extended states) correspond to the complex spectra. Such complex spectra form closed loops in the complex plane and cause the system to exhibit non-zero topological winding numbers ZGong2018 ; KZhang2020 ; PHe2020 . The Anderson transition point in energy coincides with the PT symmetry breaking LWZhou2022 ; SLonghi2019 ; TLiu2022a ; XXia2022 ; SLonghi2021 ; YLiu2021c ; YLiu2021 ; YLiu2020 ; YLiu2021a .
So far, although a lot of efforts have been made in the study of the traditional point gap with winding number equal to one SLonghi2019 ; QBZeng2020b ; KKawabata2019 ; SLonghi2019a , few papers discuss the point gap with higher-spectrum winding numbers YLiu2021 ; APadhan2024 . This work, therefore, is devoted to studying the non-Hermitian topological point gap with higher-spectrum winding numbers, the corresponding localization properties and MEs.
The main findings of this manuscript are as follows. Multiple MEs and Matryoshka doll-like higher-spectrum point gap can emerge in non-Hermitian quasiperiodic lattice system.
The rest of the paper is structured as follows. In Sec. II, we give a brief introduction to the theoretical model and the main findings. In Sec. III, through three cases, we provide the evidence of the emergent Matryoshka doll-like point gaps. By Avila’s global theory, we analytially discuss the case of infinity summation limit in Sec. IV. We summarize this manuscript in Sec. V.
II Model
Let’s start at a non-Hermitian geometric series modulated (GSM) qusiperiodic potential [see Fig. 1(a)]. The corresponding Hamiltonian reads
$H=\sum_{j=1}^{L-1}(Jc_{j}^{\dagger}c_{j+1}+\mathrm{H.c.})+\sum_{j=1}^{L}V_{j}c_{j}^{\dagger}c_{j},$ (1)
where
$V_{j}=\sum_{n=1}^{N}{\lambda}e^{-(n-1)q}e^{in(2{\pi}{\alpha}j+{\theta})},$ (2)
and $c_{j}^{\dagger}$ ($c_{j}$) denotes the fermion creation (annihilation) operator on site $j$. $J$ represents the nearest-neighboring hopping strength and $L$ is the system size. $V_{j}$ is the GSM potential with quasiperiodic strength $\lambda$. $N$ is the number of summation terms of the geometric series. The corresponding common ratio of the geometric series is $e^{-q}e^{i(2\pi\alpha j+\theta)}$, where $\alpha$ is an irrational number to garantee that the potential energy has a quasiperiodic structure and $\theta$ is the phase offset.
Figure 1: The scheme diagram of GSM lattice model (a) and Matryoshka doll-like point gaps spectra (b).
When $\theta=0$, one has $V_{-j}=V_{j}^{\ast}$ and the system is PT symmetric in rigorous periodic boundary conditions (PBCs). Without loss of generality, in numerical computation, we set $J=1$ as unit energy, $\theta=0$, $q=0.2$ and $\alpha=\lim_{m\rightarrow\infty}F_{m-1}/F_{m}=(\sqrt{5}-1)/2$ with $F_{m}$ being the $m$-th Fibonacci number. For finite size cases, we set system size $L=F_{m}$ and $\alpha=F_{m-1}/F_{m}$ to ensure rigorous PBCs.
We focus on the properties of MEs and topology point gaps. To this end, we calculate two key quantities.
One is the fractal dimension, which constitutes the core quantity describing the localization properties and MEs. The fractal dimension can be defined as YWang2020
$\Gamma_{m}=-\lim_{L\rightarrow\infty}\frac{\ln\xi_{m}}{\ln L},$ (3)
where $\xi_{m}=\sum_{j=1}^{L}|\psi_{j,m}|^{4}$ is the corresponding inverse participation ratio. $\psi_{j,m}$ denotes the amplitude for the $m$th eigenstate at the $j$th site. For the localized (extended) state, $\Gamma=0$ ($\Gamma=1$), while the critical state corresponds to $\Gamma\in(0,1)$. Note that, $\Gamma$ tending to $1(0)$ would indicate the extended (localized) state at finite size.
The other is the winding number, which represents a key quantity that reflects the topological properties and can be defined as
[…]
which describes the number of times the spectrum revolving around the base point $E_{b}$ in the complex plane ZGong2018 ; SLonghi2019 . Below, we exhibit that Matryoshka doll-like point gaps [see Fig. 1(b)] emerge in the lattice model (1).
III Matryoshka doll-like point gaps
Figure 2: (a) The phase diagram of fractal dimension ${\Gamma}$ in the $\lambda-E$ plane. Eigenenergies in the complex plane with $\lambda=0.2$ (b), $\lambda=0.7$ (c), $\lambda=1$ (d) and $\lambda=1.7$ (e). (f) Arg[det($H-E_{b}$)] versus $\theta$, where we set $E_{b}=1,~{}1,~{}1,~{}-0.95$ [marked as rhombus in (b)-(e)]. Throughout, we set $L=610$, $\alpha=377/610$ and $N=2$.
III.1 Traditional point gap with $\omega=1$
First, we consider the case of $N=1$. The quasiperiodic potential contains only the first term of the geometric series (2), i.e.,
$V_{j}=\lambda e^{(i2\pi\alpha j+\theta)}.$ (5)
Under such circumstance, the model is an integrable system. The critical point of Anderson transition is located at $\lambda_{c}=J$. The corresponding Lyapunov exponent $\gamma=\ln(\lambda/J)$ SLonghi2019 ; SLonghi2019a ; YLiu2021 . When $\lambda<J$ ($\lambda>J$), the system is extended (localized) and the eigenenergy spectrum $E(k)=2\cos(k)$ [$E(k)=\lambda e^{ik}+\lambda e^{-ik}$] with $k\in(0,2\pi]$, which corresponds to the trivial (topological) phase. In the topological phase, the maximum winding number $\omega=1$, which equals to the summation number of the GSM potential.
III.2 The point gap with $\omega=2$
Now, let’s turn to point gaps with hight winding numbers. Under the condition of $N=2$, the corresponding quasiperiodic potential reads
$V_{j}={\lambda}e^{i(2{\pi}{\alpha}j+{\theta})}+\lambda e^{-q}e^{i2(2\pi\alpha+j+\theta)}.$ (6)
In Fig. 2(a) we exhibit the fractal dimensions $\Gamma$ versus $\lambda$ and $E$. By contrast to the case of traditional point gap, the Matryoshka doll-like point gaps emerge here accompanied by MEs. In concrete terms, for the former, there is no ME, and only a clean critical point ($\lambda=J$) exists between the localized and extended states. For the latter, however, the potential of one more summation term cause the ME to be in the interval $1.20.48<\lambda<$ 1.2 [see Fig. 2(a)]. Furthermore, the localized states gradually invade the extended region until all eigenstates become localized with an increasing $\lambda$. Meanwhile, as the eigenspectrum changes from pure real to complex, the system’s PT symmetry is gradually broken [see Fig. 2(b)-(e)]. One can also see that the localized state forms two separated circles at the maximum and minimum eigenvalues, which are connected by the eigenvalues of the extended state. And when all the eigenstates are localized, the two circles fuse to form a Matryoshka doll-like structure [see Fig. 2(e)]. Fig. 2(f) exhibits the argument of the determinants $H-E_{b}$ as a function of the flux $\theta$, where the corresponding base points $E_{b}$ are marked as a rhombus in Fig. 2(b)-(e). One can find that the Arg[det($H-E_{b}$)] is always zero with $\theta$ for the extended phase ($\lambda=0.2$), corresponding to the trivial phase with winding number $\omega=0$. Under the condition of $\lambda=0.7$ and $\lambda=1$, the Arg[det($H-E_{b}$)] forms a cycle with the $\theta$, corresponding to the topological phase with winding numbers $\omega=1$. Under the condition of complete localization ($\lambda=1.7$), a topological phase with higher-order winding number $\omega=2$ emerges with an increasing $\theta$.
In other words, two stepwise phase transitions occur in the system, i.e., localization and topological phase transition. These two phase transitions have exactly identical critical points and corresponding regions of different phases. That is to say, when the system gradually changes from the extended phase, through the intermediate phase, and finally into the localized phase, the topological properties of the system also gradually change from the trival phase, through the topological phase of $\omega=1$, and finally into the topological phase of $\omega=2$. Meanwhile, the symmetry of the system will start with PT symmetry, then enter a partial PT symmetry broken region, and finally become a state where PT symmetry is completely broken.
Figure 3: (a) The phase diagram of fractal dimension ${\Gamma}$ in the $\lambda-E$ plane. Eigenenergies in the complex plane with $\lambda=0.8$ (b), $\lambda=1.2$ (c), $\lambda=1.9$ (d). (f) Arg[det($H-E_{b}$)] versus $\theta$, where we set $E_{b}=0,~{}0,~{}-1.85$ [marked as rhombus in (b)-(d)]. The other parameters $L=610$, $\alpha=377/610$ and $N=3$.
III.3 The point gap with $\omega=3$
For the case of $N=3$, one more term is added into the summation of GSM potential, which can be written as
$V_{j}={\lambda}e^{i(2{\pi}{\alpha}j+{\theta})}+{\lambda}e^{-q}e^{2i(2{\pi}{\alpha}j+{\theta})}+{\lambda}e^{-2q}e^{3i(2{\pi}{\alpha}j+{\theta})}.$ (7)
The corresponding fractal dimension $\Gamma$ of all eigenstates as a function of $\lambda$ is plotted in the following Fig. 3(a). One can find that there are more MEs and more circles of point gap spectrum [see Fig. 3(b)-(d)]. Furthermore, the point gap’s circles emerge one by one in the complex plane. To be more specific, two circles will appear first at either end of the real energy, and then the third circle structure appears from the middle of the energy spectrum [see Fig. 3(a)]. These circle structures with complex values correspond to localized eigenstates, which are connected together in the complex plane by the pure real eigenspectrum of extended states. Later, with the increase of disorder strength $\lambda$, the three discrete circles will become increasingly larger. Finally, they merge together to form a three-layer Matryoshka doll-like point gap [see Fig. 3(b)-(d)]. The corresponding winding numbers will gradually complete the transformation of $0\rightarrow 1\rightarrow 2\rightarrow 3$. Again, accompanied by topological and localization phase transitions, the PT symmetry gradually breaks. The corresponding arguments confirms the conclusion as shown in Fig. 3(e).
Combining the above results of $N=2$ and $N=3$, one can find that the winding number, the circle number of point gaps and the number of MEs are all dominated by the potential energy structure, i.e. the number of summation terms $N$. In concrete terms, the number of summation terms $N$ corresponds to the winding number $\omega=N$, the circle number of point gaps is equal to $N$, and the number of MEs is equal to $2(N-1)$.
IV Theoretical analysis in infinite limit ($N={\infty}$)
Now, let’s consider the limit case of $N=\infty$. Since the common ratio satisfies $|e^{-q}e^{i(2{\pi}{\alpha}j+{\theta})}|\in(0,~{}1]$, the corresponding geometric series converges. Then, by using the geometric series’ summation formula, one can obtain
$V_{j}=\frac{{\lambda}e^{i(2{\pi}{\alpha}j+{\theta})}}{1-e^{-q}e^{i(2{\pi}{\alpha}j+{\theta})}}.$ (8)
Note that, the expression Eq. (8) is a non-Hermitian version Ganeshan-Pixley-Das Sarma potential SGaneshan2015 , and the corresponding MEs can likewise be solved exactly by the Avila global theory.
Under the condition of $N=\infty$, the eigenequation for Hamiltonian (1) is
$E\psi_{j}=\psi_{j+1}+\psi_{j-1}+V_{j}\psi_{j},$ (9)
and the corresponding transfer matrix is
$T_{j}=\begin{bmatrix}E-V_{j}&-1\\ 1&0\end{bmatrix}=A_{j}B_{j},$ (10)
where
$A_{j}=\frac{1}{1-e^{-q}e^{i(2{\pi}{\alpha}j+{\theta})}},$ (11)
$B_{j}=\begin{bmatrix}E-(Ee^{-q}+\lambda)e^{i(2\pi\alpha j+\theta)}&-1+e^{-q}e^{i(2\pi\alpha j+\theta)}\\ 1-e^{-q}e^{i(2\pi\alpha j+\theta)}&0\end{bmatrix}.$
With the help of the transfer matrix, one can calculate the Lyapunov exponent (LE), i.e.,
$\gamma=\lim_{L\rightarrow\infty}\frac{1}{L}\ln\left\|{\textstyle\prod_{j=1}^{L}}T_{j}\right\|,$ (12)
[…]
$\gamma_{A}(E)=\lim_{L\rightarrow\infty}\frac{1}{L}\ln\prod_{j=1}^{L}\frac{1}{\left|1-e^{-q}e^{i(2\pi\alpha j+\theta)}\right|}=\left\{\begin{matrix}0,&q\geq 0,\\ -q,&q<0.\end{matrix}\right.$ (13)
and $\gamma_{B}(E)$ can be calculated by Avila global theroy AAvila2015 . We insert an additional imaginary phase $\theta\rightarrow\theta+i\epsilon$. Although $T_{j}\in SL(2,\mathbb{C})$ is not an even function for $\epsilon$, we can handle it by complexification as well YCZhang2022 ; YWang2021 ; YWang2020a ; YWang2021a ; AAvila2015 . Then, by letting $\epsilon\rightarrow-\infty$, one can get
$B_{j,\epsilon}=e^{-q}e^{\epsilon}e^{i(2\pi\alpha j+\theta)}\begin{bmatrix}-(E+\lambda e^{q})&1\\ -1&0\end{bmatrix}+\mathcal{O}(1).$ (14)
Then, we obtain $\gamma_{B,\epsilon}=\gamma_{B,0}+\epsilon+\mathcal{O}(1)$, where $\gamma_{B,0}=\ln\left|\frac{|E+\lambda e^{q}|+\sqrt{|E+\lambda e^{q}|-4}}{2}\right|-q$. Combined with $\gamma_{A}$, we obtain $\gamma_{\epsilon}=\gamma_{A}+\gamma_{B,0}+\epsilon+\mathcal{O}(1)$. Note that, $\gamma_{\epsilon}$ is a convex, piecewise linear function of $\epsilon$ with their slopes being integers. If the energy belongs to the spectrum, one can obtain the LE for $q>0$, i.e.,
$\gamma=\ln\left|\frac{|E+\lambda e^{q}|+\sqrt{|E+\lambda e^{q}|-4}}{2}\right|-q.$ (15)
The ciritical energy of MEs satisfies $\gamma=0$. Then, one can obtain
$E_{c}=(1-\lambda)e^{q}+e^{-q}.$ (16)
The numerical fractal dimension $\Gamma$ as a function of $\lambda$ are plotted in Fig. 4(a), where the red dashed line is the exact ME $E_{c}$ from the expression Eq. (16). The numerical results are in good agreement with analytical expression.
From the Fig. 4(a), as the quasiperiodic strength $\lambda$ increases, the first phase transition occurs at $E=2$, and the last phase transition occurs at $E=-2$ ($\lambda_{c}\approx 3.5$). After that, the system enters the localized phase completely. For the case of $-2<E<2$, the critical points are in the interval $3.50<\lambda_{c}<$ 3.5.
Furthermore, we can fix the base points $E_{b}=\pm 2$ and compute the winding number $\omega$ as a function of $\lambda$ to characterize the appearance of MEs and the complete localization transition, respectively [see the insets of Fig. 4(b)]. One can see that when the MEs appear, the $\omega$ for the base $E_{b}=-2$ transforms from 0 to 1 and the system has a topological point gap. When $\omega$ for the base point $E_{b}=\pm 2$ are both non-zero, the system enters the localized phase completely. Specifically, we show in the inset the fractal dimension $\Gamma$ in the complex plane under $\lambda=1$ (coexistence of extended and localized states) and $\lambda=4$ (localized phase), which agree with our previous analysis well. Therefore, for the case of $N=\infty$, the localized region encroaches on the extended region from the maximum real energy, and there is always only one topological gap in the system. The corresponding maximum winding number $\omega=1$ (see Appendix B for details about phase diagrams versus $N$).
Figure 4: (a) The phase diagram of fractal dimension ${\Gamma}$ in the $\lambda-E$ plane, red dashed line denotes the analytical ME. (b) The winding number versus $\lambda$ for $E_{b}=-2$ (red line) and $E_{b}=2$ (blue line). The insets exhibit point gaps in the complex plane, where the base point $E_{b}=-2,~{}2$ [the rhombus in the insets]. Parameters $L=610$, $\alpha=377/610$ and $N=\infty$.
V Conclusion
In summary, we study localization and topological phase transitions in a GSM quasiperiodic lattice. The results reveal that multiple MEs and point gaps with high winding number can emerge under the condition of summation terms’ number $N>1$. The system’s PT symmetry is damaged synchronously with the phase transition. Specifically, for the case of finite summation number $N$, the circle number of topological point gaps can at most be $N$. As the eigenstates continuously transform into localized states, the different point gaps gradually merge, and the maximum winding number gradually increases from $\omega=0$ to $\omega=N$. For the limit case of $N=\infty$, however, Avila’s global theory can give the exact ME’s expression. Analytical and numerical results consistently show that the corresponding MEs (topological point gaps) will merge together, resulting in a winding number $\omega=1$. Furthermore, the main results confirm that the properties of MEs and point gap structures are consistent with the behavior of topological boundaries and PT symmetry breaking.
VI Acknowledgements
We thank R.-J. Chen, X.-D. Hu and W.-L. Li for their insightful suggestions. This work was supported by the National Key Research and Development Program of China (Grant No.2022YFA1405300), and the Guangdong Basic and Applied Basic Research Foundation (Grants No.2021A1515012350).
Figure 5: The detialed evolution of point gap and the corresponding winding number with repsect to $\lambda$ under the condition of $N=2$. The qusiperiodic strength are marked. The other parameters $L=610$, $\alpha=377/610$ and $q=0.2$.
Figure 6: The detialed evolution of point gap and the corresponding winding number with repsect to $\lambda$ under the condition of $N=4$. The qusiperiodic strength are marked. The other parameters $L=610$, $\alpha=377/610$ and $q=0.05$.
Appendix A The evolution of point gap with respect to $\lambda$
To exhibit the Matryoshka doll-like point gaps and the process of finding the base point $E_{b}$, we provide two examples, i.e., the case of $N=2$ and $N=4$.
First, let’s discuss the case of $N=2$. It is not difficult to find from Fig. 5 that with the increase of quasiperiodic strength, some states with the largest eigenenergy in the system undergo localization phase transitions first. Then, as $\lambda$ continues to increase, some eigenstates with the lowest energy begin to enter the localized state. Then, the number of localized states will continue to increase with an increasing $\lambda$, and eventually all extended states will tranform into localized states. Then, the two points will merge together to form a two-layer Matryoshka doll-like structure. After that, by further increasing $\lambda$, the size of the ring become even larger, and the two-layer circile structure will always exist.
Now we turn to the case of $N=4$. With the increase of quasiperiodic strength $\lambda$, the pace of phase transition is simular to the former case, i.e., the states in the two ends of eigenenergy localized first, and the states in the middle will follow suit. Besides, unlike the $N=2$ case, there are more point gaps with higher-order winding number. Specifically, there are point gaps with winding number $\omega=1$ (e.g., subfigures of $\lambda=0.4$), $\omega=2$ (e.g., subfigures of $\lambda=1.0$), point gaps of $\omega=3$ (e.g., subfigures of $\lambda=1.4$), $\omega=4$ (e.g., subfigures of $\lambda=1.6$) [see Fig. 6].
Appendix B Discussions on phase diagram of different $N$
In fact, while the number of MEs increases with an increasing $N$, the interval between MEs decreases versus $N$. Then, for the case of $N\rightarrow\infty$, the interval between two nearest-neighboring MEs approaches 00, and finally merges into one ME [see Fig. 7].
Figure 7: $\Gamma$ phase diagram on $\lambda-E$ plane with different $N$. Parameters $L=610$, $\alpha=377/610$ and $q=0.2$.
The relationship between the ME and the parameter $N$ can be summarized as follows. Under the condition of $N=1$, the system exhibits a traditional topological point gap, in which there is no ME and only a standard critical point, namely $\lambda_{c}=J$. With the increase of parameter $N$, multiple MEs gradually appear in the system, and the number of MEs is $2(N-1)$, which is consistent with the previous prediction. Finally, as $N$ increases, the exponential decay term at large $N$ makes the MEs all merge together, while only one ME can be observed on the phase diagram. One can obtain the same conclusion from the point of view of topological point gap analysis, i.e., although the increase of point gaps’s circle promotes a higher winding number with an increasing $N$, the maximum winding number decreases to $1$ due to the merging of point gaps’ circles.